OpenStudy (amistre64):

integrate; 27x sec^2 (3x) dx

OpenStudy (anonymous):

wolframalpha?

OpenStudy (amistre64):

\[27x\frac{tan(3x)}{3}-27(??)+C\]is what i got so far :)

OpenStudy (amistre64):

why wolfram? i dont care about an answer, its the process i enjoy

OpenStudy (anonymous):

because if u use hashtag evaluate it gives the process, just like mathematica 8

OpenStudy (amistre64):

if I can see how to integrate tan(3x) id be doing good :)

OpenStudy (amistre64):

cool, ill have to try to remember the hashtag trick then

OpenStudy (amistre64):

in the mean time, im practiceing our upcoming int by parts for calcII

OpenStudy (amistre64):

maybe u = 27 and dv = x sec^2(3x) ?

OpenStudy (amistre64):

i know, that was pointless :)

OpenStudy (amistre64):

sec^2 = tan^2 + 1\[\] 27x tan^2(3x) +27x ... but then what to do with the tan^2(3x)\[\] i cant recall to well

OpenStudy (zarkon):

integration by parts

OpenStudy (zarkon):

u=27x

OpenStudy (amistre64):

ln(cos(3x)) somehow

OpenStudy (amistre64):

i got that much :) its that sec^2(3x) part i got a brainurysm on

OpenStudy (zarkon):

\[\frac{\tan(3x)}{3}\]

OpenStudy (amistre64):

right, then that suits up to ln(cos(3x)) somehow :)

OpenStudy (amistre64):

or something akin to it

OpenStudy (zarkon):

\[\int\tan(x)dx=\int\frac{sin(x)}{\cos(x)}dx=\int\frac{-1}{u}du=-\ln(|u|)+c=-\ln|cos(x)|+c\]

OpenStudy (amistre64):

... yep, thats what I was brain blocked on lol, thnx :)

OpenStudy (zarkon):

np